# Green’s matching: an efficient approach to parameter estimation in complex dynamic systems

### v1

@@ 2.4 Optimization scheme for Green’s matching
and its derivative of $\boldsymbol{\beta}$ is $-2\sum_{i=1}^{p}\big{[}\boldsymbol{\Psi}_{i}\left\{\nabla\boldsymbol{\nu}_{i}(\boldsymbol{\beta})\right\}+\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\big{]}^{T}\big{\{}\boldsymbol{Z}_{i}-\boldsymbol{\Psi}_{i}\boldsymbol{\nu}_{i}(\boldsymbol{\beta})-\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\big{\}}$, where $\nabla$ is the gradient operator. Since $\nabla\boldsymbol{\nu}_{i}(\boldsymbol{\beta})=-(\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Psi}_{i})^{-1}\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Phi}_{i}\left\{\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\right\}$, we can specify $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ and $\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ to calculate the above gradient. It’s worth nothing that (17) can be considered a nonlinear least-squares problem. Therefore, it can be efficiently solved by many existing Newton-type methods given the above gradients; see Nocedal and Wright (1999) for more details of these methods.
@@ 3.1 Statistical properties of Green’s matching
The observed time points $\{t_{j};j=1,\cdots,n\}$ are the i.i.d. random variables with density function $f_{T}(t)$, where $f_{T}(t)$ is bounded away from 00 and continuous on $[0,C]$.
@@
In the following, we denote the $u^{th}$ component of $\frac{\partial f_{i}(\boldsymbol{X};\boldsymbol{\beta})}{\partial\boldsymbol{X}}$ as $\frac{\partial f_{i}(\boldsymbol{X};\boldsymbol{\beta})}{\partial X_{u}}$. The notations $\frac{\partial f_{i}(\boldsymbol{X}(s);\boldsymbol{\beta}_{0})}{\partial\boldsymbol{\beta}}$ and $\frac{\partial f_{i}(\boldsymbol{X}(s);\boldsymbol{\beta}_{0})}{\partial X_{u}}$ are also simplified as $\frac{\partial f_{i}(s)}{\partial\boldsymbol{\beta}}$ and $\frac{\partial f_{i}(s)}{\partial X_{u}}$, respectively. Besides, we define $\boldsymbol{d}_{i}(t)=(\int_{0}^{1}G_{i}^{K}(t,s)\frac{\partial f_{i}(s)}{\partial\boldsymbol{\beta}^{T}}\mathrm{d}s,\ \boldsymbol{e}_{i}^{T}\otimes\{\boldsymbol{\psi}_{i}(t)\}^{T})^{T}$, where $\boldsymbol{e}_{i}\in\mathbb{R}^{p}$ satisfies $[\boldsymbol{e}_{i}]_{i^{\prime}}=\mathbb{I}(i=i^{\prime})$ and $\otimes$ is the Kronecker product.
@@ 3.2 Convergence rates of three methods
(b) Under Assumptions A.1-A.3, A.5, B.1-B.4, B.6, B.7, and $h_{k}=o(1)$ such that $n^{-1}\ln n/h_{k}^{2{k}+1}\to 0$ as $n\to\infty$ for ${k}\leq K_{0}$, where $K_{0}=K$ or $K-1$, then
@@
for $K_{0}=K$ or $K-1$.
Combining with Theorems 2 and 3, the main difference of the asymptotic behaviors between Green’s matching and gradient matching lies in the presence of $\boldsymbol{W}_{n,K_{0}}$, where the convergence rate gets slow as $K_{0}$ increases. Based on this, the convergence orders of the estimation biases increase in the order of Green’s matching, order $(K-1)$ gradient matching, and order $K$ gradient matching, where the corresponding used pre-smoothing is increased in turn. Typically, the pre-smoothing step would enlarge the total bias of the resulting estimator by ignoring prior information of the underlying model. Therefore, it can be expected that the larger amount of pre-smoothing used in the first step, the greater bias in the second-step estimation.
@@
when $h_{0},\ h_{1},\cdots,\ h_{K-1}$ are of the order $n^{-\frac{1}{K+2}}$. If $K=1$, these results are consistent with Brunel (2008), Gugushvili and Klaassen (2012), and Dattner and Klaassen (2015) for first-order systems. However, as $K$ becomes larger, the biases would dominate the convergence rates, leading to the order $K$ or $(K-1)$ gradient matching may not achieve $\sqrt{n}$-consistency when $K\geq 2$ or $K\geq 3$, respectively. Meanwhile, Green’s matching can always obtain the $\sqrt{n}$-consistency.
@@ 4.1 Comparisons of two-step methods
For the above equations, we simulate their dynamic curves using the corresponding true values of parameters and the initial conditions mentioned above. The Fig. 1 in Supplementary Materials shows the simulated dynamic curves. Next, we generate $Y_{i}(t_{j})$ from a normal distribution with mean $X_{i}(t_{j})$ and variance $\sigma_{i}^{2}$, where $\sigma_{i}^{2}=\gamma^{2}\int_{0}^{C}X_{i}^{2}(t)\ \mathrm{d}t/C$ with $\gamma$ controlling the contamination level. It’s worth noting that Wu et al. (2014) and Zhang et al. (2020) inferred the gene regulatory network and the oscillatory dynamic directional model by gradient matching. Here, we further examine the estimation behaviors of Green’s matching for these dynamic systems.
@@
We use the order $K$ and $(K-1)$ gradient matching and Green’s matching to fit the observed data. Since Green’s matching is equivalent to order $(K-1)$ gradient matching in the gene regulatory network, we only compare the estimation behaviors of Green’s matching and order $K$ gradient matching in this case.
@@
The results with $G=100$ and different values of $n$s and $\gamma$s are given in Table 1. We find that Green’s matching outperforms the gradient matching at different orders for all the test models under different settings. In particular, the gradient matching approaches exhibit poor performances for second-order systems in cases with a high contamination level $\gamma$, where the pre-smoothed derivatives are usually inaccurate. In contrast, Green’s matching maintains satisfactory performances in such scenarios. These results suggest that the pre-smoothed derivatives significantly impact the estimation accuracy of the two-step procedures. In this regard, Green’s matching removes all the pre-smoothed derivatives for the second-step estimation, thereby obtaining a better estimation performance than the other two methods.
Figure 2: RBIAS and RSD of three methods with different sample sizes $n$ and $\gamma$s (sub-title), where RSDs are colored in gray, and RBIASs of order 2 / 1 gradient matching and Green’s matching are colored in blue, orange, and red, respectively.
@@
where $\bar{\beta}_{l}=\frac{1}{G}\sum_{g=1}^{G}{\beta}^{(g)}_{l}$. RBIAS and RSD are two additional measures to assess the corresponding estimators’ relative estimation biases and variances. In Fig. 2, we present the RBIASs and RSDs of the three methods for all models. We find that the estimation biases of gradient matching always dominate their estimation errors, whereas those of Green’s matching are nearly balanced with the corresponding RSDs. These results are consistent with Theorem 3. For a two-step estimator, the pre-smoothing process tends to enlarge the total estimation bias, leading to a poor convergence and/or finite-sample behavior for parameter estimation. However, these estimation biases can be alleviated by using Green’s matching, which results in superior performances for parameter estimation in general-order systems.
For subsequent analysis, we only focus on the oscillatory dynamic directional model with $n=250$ and $\gamma=7\%$. To illustrate the ability to capture data patterns, we use a numerical approximation to reconstruct the trajectories $X_{1}(\cdot)$, $DX_{1}(\cdot)$, and $D^{2}X_{1}(\cdot)$ based on the pre-smoothed initial conditions and the estimated parameters obtained from the three methods. The results are shown in Fig. 3 A. For these cases, the order 2 gradient matching introduces $\widehat{D^{2}X}_{1}(\cdot)$ to the second-step parameter estimation, even though $D^{2}X_{1}(t)$ is discontinuous at $t=2$ or $t=3$. We observe that the reconstructed curves by order 2 gradient matching poorly fit their true shapes. In contrast, those by the other two methods appropriately fit the underlying curves, as they remove the pre-smoothed $D^{2}X_{1}(t)$ from the parameter estimation. These results suggest that incorporating pre-smoothing of non-smooth derivatives into the two-step methods may result in significant errors in parameter estimation.
@@
Furthermore, we calculate the 95% confidence intervals for the parameters $a_{1}$, $b_{1}$, $c_{1}$, and $d_{1}$ in the oscillatory dynamic directional model, utilizing the asymptotic normality of Green’s matching in Theorem 2. We also employ a similar strategy to construct the 95% confidence intervals by gradient matching of different orders. These 95% confidence intervals are shown in Fig. 3. B. In our simulation, nearly 95% (94%, 92%, 96%, 88% for $a_{1}$, $b_{1}$, $c_{1}$, and $d_{1}$, respectively) of the confidence intervals constructed by Green’s matching cover the true parameters. While the confidence intervals constructed by gradient matching are significantly biased with incorrect covering proportions (correspondingly, 39%, 42%, 47%, and 7% in order $K$ gradient matching, and 0%, 0%, 1%, and 0% in order $(K-1)$ gradient matching).
@@ 5 Equation discovery of a nonlinear pendulum
We generate 100 replications of simulated data for the nonlinear pendulum by setting $n=50$ and $\gamma=5\%$, in which the initial values $(X(0),DX(0))$ are randomly sampled from the uniform distribution on $[-0.5,0.5]$. After that, we respectively apply gradient matching and Green’s matching for discovering equations from the observed data. It’s worth noting that the discovered equation determines $(DX(t),D^{2}X(t))$ given $(X(t),DX(t))$, i.e., determining a vector field from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$. To compare different methods, we present the true and the estimated vector fields for the nonlinear pendulum in Figure 4. From these results, we observe that the estimated vector fields by Green’s matching are nearly the same as the true case, whereas those obtained by gradient matching exhibit significant biases outside the central region. These biases may arise from the approximation of high-order derivatives in the gradient matching, prohibiting its effectiveness in discovering equations from discrete noisy data. In contrast, Green’s matching demonstrates its superiority in equation discovery by only approximating $X(\cdot)$ for the estimation.
@@ 6.2 Extensions for complex tasks
Acknowledgments
Xueqin Wang’s research is partially supported by the National Key R&D Program of China (No. 2022YFA1003803) and the National Natural Science Foundation of China (No. 12231017, 72171216). Fang Yao’s research is partially supported by the National Key R&D Program of China (No. 2022YFA1003801, 2020YFE0204200), the National Natural Science Foundation of China (No. 12292981, 11931001, 11871080), the LMAM and the LMEQF.

### v2

@@ 2.4 Optimization scheme for Green’s matching
and its derivative of $\boldsymbol{\beta}$ is $2\sum_{i=1}^{p}\big[\boldsymbol{\Psi}_{i}\left\{\nabla\boldsymbol{\nu}_{i}(\boldsymbol{\beta})\right\}+\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\big]^{T}\big\{\boldsymbol{Z}_{i}-\boldsymbol{\Psi}_{i}\boldsymbol{\nu}_{i}(\boldsymbol{\beta})-\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\big\}$, where $\nabla$ is the gradient operator. Since $\nabla\boldsymbol{\nu}_{i}(\boldsymbol{\beta})=-(\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Psi}_{i})^{-1}\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Phi}_{i}\left\{\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\right\}$, we can specify $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ and $\nabla\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ to calculate the above gradient. It’s worth nothing that (17) can be considered a nonlinear least-squares problem. Therefore, it can be efficiently solved by many existing Newton-type methods given the above gradients; see Nocedal and Wright (1999) for more details of these methods.
@@ 3.1 Statistical properties of Green’s matching
The observed time points $\{t_{j};j=1,\cdots,n\}$ are the i.i.d. random variables with density function $f_{T}(t)$, where $f_{T}(t)$ is bounded away from 0 and continuous on $[0,C]$.
@@
In the following, we denote the $u^{th}$ component of $\frac{\partial f_{i}(\boldsymbol{X};\boldsymbol{\beta})}{\partial\boldsymbol{X}}$ as $\frac{\partial f_{i}(\boldsymbol{X};\boldsymbol{\beta})}{\partial X_{u}}$. The notations $\frac{\partial f_{i}(\boldsymbol{X}(s);\boldsymbol{\beta}_{0})}{\partial\boldsymbol{\beta}}$ and $\frac{\partial f_{i}(\boldsymbol{X}(s);\boldsymbol{\beta}_{0})}{\partial X_{u}}$ are also simplified as $\frac{\partial f_{i}(s)}{\partial\boldsymbol{\beta}}$ and $\frac{\partial f_{i}(s)}{\partial X_{u}}$, respectively. Besides, we define $\boldsymbol{d}_{i}(t)=(\int_{0}^{1}G_{i}^{K}(t,s)\frac{\partial f_{i}(s)}{\partial\boldsymbol{\beta}^{T}}\mathrm{d}s,\ \boldsymbol{e}_{i}^{T}\otimes\{\boldsymbol{\psi}_{i}(t)\}^{T})^{T}$, where $\boldsymbol{e}_{i}\in\mathbb{R}^{p}$ satisfies $\boldsymbol{e}_{i}]_{i^{\prime}}=\mathbb{I}(i=i^{\prime})$ and $\otimes$ is the Kronecker product.
@@ 3.2 Convergence rates of three methods
(b) Under Assumptions A.1-A.3, A.5, B.1-B.4, B.6, B.7, and $h_{k}=o(1)$ such that $n^{-1}\ln n/h_{k}^{2{k}+1}\to 0$ as $n\to\infty$ for ${k}\leq K_{0}$, where $K_{0}=K$ or K1K-1, then
@@
for $K_{0}=K$ or K1K-1.
Combining with Theorems 2 and 3, the main difference of the asymptotic behaviors between Green’s matching and gradient matching lies in the presence of $\boldsymbol{W}_{n,K_{0}}$, where the convergence rate gets slow as $K_{0}$ increases. Based on this, the convergence orders of the estimation biases increase in the order of Green’s matching, order (K1)(K-1) gradient matching, and order $K$ gradient matching, where the corresponding used pre-smoothing is increased in turn. Typically, the pre-smoothing step would enlarge the total bias of the resulting estimator by ignoring prior information of the underlying model. Therefore, it can be expected that the larger amount of pre-smoothing used in the first step, the greater bias in the second-step estimation.
@@
when $h_{0},\ h_{1},\cdots,\ h_{K-1}$ are of the order $n^{-\frac{1}{K+2}}$. If $K=1$, these results are consistent with Brunel (2008), Gugushvili and Klaassen (2012), and Dattner and Klaassen (2015) for first-order systems. However, as $K$ becomes larger, the biases would dominate the convergence rates, leading to the order $K$ or (K1)(K-1) gradient matching may not achieve $\sqrt{n}$-consistency when $K\geq 2$ or $K\geq 3$, respectively. Meanwhile, Green’s matching can always obtain the $\sqrt{n}$-consistency.
@@ 4.1 Comparisons of two-step methods
For the above equations, we simulate their dynamic curves using the corresponding true values of parameters and the initial conditions mentioned above. The Figure 1 in Supplementary Materials shows the simulated dynamic curves. Next, we generate $Y_{i}(t_{j})$ from a normal distribution with mean $X_{i}(t_{j})$ and variance $\sigma_{i}^{2}$, where $\sigma_{i}^{2}=\gamma^{2}\int_{0}^{C}X_{i}^{2}(t)\ \mathrm{d}t/C$ with $\gamma$ controlling the contamination level. It’s worth noting that Wu et al. (2014) and Zhang et al. (2020) inferred the gene regulatory network and the oscillatory dynamic directional model by gradient matching. Here, we further examine the estimation behaviors of Green’s matching for these dynamic systems.
@@
We use the order $K$ and (K1)(K-1) gradient matching and Green’s matching to fit the observed data. Since Green’s matching is equivalent to order (K1)(K-1) gradient matching in the gene regulatory network, we only compare the estimation behaviors of Green’s matching and order $K$ gradient matching in this case.
@@
[…]
@@
where $\bar{\beta}_{l}=\frac{1}{G}\sum_{g=1}^{G}{\beta}^{(g)}_{l}$. RBIAS and RSD are two additional measures to assess the corresponding estimators’ relative estimation biases and variances. In Figure 2, we present the RBIASs and RSDs of the three methods for all models. We find that the estimation biases of gradient matching always dominate their estimation errors, whereas those of Green’s matching are nearly balanced with the corresponding RSDs. These results are consistent with Theorem 3. For a two-step estimator, the pre-smoothing process tends to enlarge the total estimation bias, leading to a poor convergence and/or finite-sample behavior for parameter estimation. However, these estimation biases can be alleviated by using Green’s matching, which results in superior performances for parameter estimation in general-order systems.
For subsequent analysis, we only focus on the oscillatory dynamic directional model with $n=250$ and $\gamma=7\%$. To illustrate the ability to capture data patterns, we use a numerical approximation to reconstruct the trajectories $X_{1}(\cdot)$, $DX_{1}(\cdot)$, and $D^{2}X_{1}(\cdot)$ based on the pre-smoothed initial conditions and the estimated parameters obtained from the three methods. The results are shown in Figure 3 A. For these cases, the order 2 gradient matching introduces $\widehat{D^{2}X}_{1}(\cdot)$ to the second-step parameter estimation, even though $D^{2}X_{1}(t)$ is discontinuous at $t=2$ or $t=3$. We observe that the reconstructed curves by order 2 gradient matching poorly fit their true shapes. In contrast, those by the other two methods appropriately fit the underlying curves, as they remove the pre-smoothed $D^{2}X_{1}(t)$ from the parameter estimation. These results suggest that incorporating pre-smoothing of non-smooth derivatives into the two-step methods may result in significant errors in parameter estimation.
@@
Furthermore, we calculate the 95% confidence intervals for the parameters $a_{1}$, $b_{1}$, $c_{1}$, and $d_{1}$ in the oscillatory dynamic directional model, utilizing the asymptotic normality of Green’s matching in Theorem 2. We also employ a similar strategy to construct the 95% confidence intervals by gradient matching of different orders. These 95% confidence intervals are shown in Figure 3. B. In our simulation, nearly 95% (94%, 92%, 96%, 88% for $a_{1}$, $b_{1}$, $c_{1}$, and $d_{1}$, respectively) of the confidence intervals constructed by Green’s matching cover the true parameters. While the confidence intervals constructed by gradient matching are significantly biased with incorrect covering proportions (correspondingly, 39%, 42%, 47%, and 7% in order $K$ gradient matching, and 0%, 0%, 1%, and 0% in order (K1)(K-1) gradient matching).
@@ 5 Equation discovery of a nonlinear pendulum
We generate 100 replications of simulated data for the nonlinear pendulum by setting $n=50$ and $\gamma=5\%$, in which the initial values $(X(0),DX(0))$ are randomly sampled from the uniform distribution on [0.5,0.5][-0.5,0.5]. After that, we respectively apply gradient matching and Green’s matching for discovering equations from the observed data. It’s worth noting that the discovered equation determines $(DX(t),D^{2}X(t))$ given $(X(t),DX(t))$, i.e., determining a vector field from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$. To compare different methods, we present the true and the estimated vector fields for the nonlinear pendulum in Figure 4. From these results, we observe that the estimated vector fields by Green’s matching are nearly the same as the true case, whereas those obtained by gradient matching exhibit significant biases outside the central region. These biases may arise from the approximation of high-order derivatives in the gradient matching, prohibiting its effectiveness in discovering equations from discrete noisy data. In contrast, Green’s matching demonstrates its superiority in equation discovery by only approximating $X(\cdot)$ for the estimation.